OpenStudy (anonymous):

Solve and express the solution set in simplest form. 6x-1=3/1 and 6x-1over 5

OpenStudy (anonymous):

6x-1/5 =3/1 6x=3+1/5 6x=16 1/5 x=16/30 x=8/15

OpenStudy (highlyeducated87):

Yep it's right were u asking just to be sure

OpenStudy (highlyeducated87):

sorry it took me so long i was doing something else

OpenStudy (anonymous):

Yeah.

OpenStudy (highlyeducated87):

Ok ur correct